# HIRO: Hierarchical Information Retrieval Optimization

### v1

@@ Introduction
RAPTOR with HIRO Querying 1864.29344498
Table 1: Average Context Length (in tokens) Returned With and Without HIRO on the NarrativeQA Dataset.
Model ROUGE-L F1 BLEU-1 BLEU-4 METEOR
RAPTOR with Collapsed Tree Querying 0.096542025 0.165350602 0.084131409 0.125686168
RAPTOR with HIRO Querying 0.120835776 0.173288844 0.089863157 0.134032758
Table 2: NarrativeQA Dataset Performance Absolute Comparison. Absolute performance comparison of RAPTOR with conventional querying and HIRO querying on the NarrativeQA dataset using SBERT, with GPT-3.5 Turbo as the language model. HIRO outperforms the existing querying mechanisms across all metrics.
Model ROUGE-L F1 BLEU-1 BLEU-4 METEOR
RAPTOR with Collapsed Tree Querying 0.012579121 0.021544661 0.010962057 0.01637651
RAPTOR with HIRO Querying 0.016045879 0.023011163 0.011933 0.017798317
Table 3: NarrativeQA Dataset Efficiency-Adjusted Performance Scores. Performance scores of RAPTOR with conventional querying and HIRO querying on the NarrativeQA dataset, using SBERT and GPT-3.5 Turbo as the language model. Scores are adjusted by the logarithm of context length (in tokens) to account for efficiency. HIRO querying demonstrates superior performance across all efficiency-adjusted metrics compared to conventional querying
@@
Table 4: Absolute Performance Metrics on the QuALITY Dataset. Performance comparison of RAPTOR with and without HIRO Querying on the QuALITY Dataset.
@@
Figure 5: Absolute Impact of HIRO on the NarrativeQA Dataset. Comparison of performance metrics between top-$k$ Collapsed Tree Querying and HIRO Querying with RAPTOR. The graph presents absolute scores for the ROUGE, BLEU-1, BLEU-4, and METEOR metrics.
Figure 6: Efficiency Analysis of HIRO on the NarrativeQA Dataset. Efficiency-adjusted performance of RAPTOR with HIRO Querying compared to Collapsed Tree Querying. HIRO Querying not only improves absolute performance but also demonstrates significant efficiency gains
@@
Figure 7: (a) Confusion Matrix for RAPTOR with Collapsed Tree Querying on the QuALITY dataset; (b) Confusion Matrix for RAPTOR with HIRO Querying on the QuALITY dataset.
Methods
Our approach introduces Recursive Similarity Score Calculation and Branch Pruning to optimize query responses through dynamic retrieval processes tailored to the specifics of each query. This method is grounded in the utilization of two hyperparameters: the Selection Threshold ($S$) and the Delta Threshold ($\Delta$), which together orchestrate a dynamic retrieval process tailored to the query’s intricacies.
Selection Threshold ($S$) serves as an initial filter, identifying document graphs for further exploration based on similarity scores between the query and parent nodes. Only graphs with parent nodes surpassing this threshold proceed, ensuring focus on potentially relevant content.
@@ Methods
10:           $local\_context.\text{extend}(\text{evaluateChildren}(q,[node],$ $\quad\quad S,\Delta))$
@@
Delta Threshold ($\Delta$) refines this by evaluating each child node within accepted graphs. Branches with similarity below $S$ are pruned, eliminating less relevant data. For remaining child nodes, a comparison of their similarity score with their parent’s determines further action. If the difference exceeds $\Delta$, indicating heightened relevance, the process recursively continues with the child node; otherwise, the parent node is retained for context.
Below, we outline the step wise functioning of this method:
  1. 1.
    Start at the root layer of the hierarchical data structure, and compute the similarity between query embedding and the embeddings of all nodes present at this initial layer
  2. 2.
    Earmark the nodes whose similarity scores surpass the predefined Selection Threshold ($S$) for further exploration
  3. 3.
    For each earmarked node, delve into its child nodes, calculating the similarity between the query embedding and each of these subsequent nodes
  4. 4.
    Compare the difference between the similarity of child node and parent node with the Delta Threshold ($\Delta$) to evaluate whether the similarity score of a child node sufficiently exceeds that of its parent node; if so, this child node is marked for further recursive evaluation. Otherwise the parent node is added to context and the remaining branch is pruned
  5. 5.
    Terminate the search if the current node is a leaf node and sufficiently exceeds the delta threshold from its parent, or when no further nodes meet the threshold criteria
  6. 6.
    The selected nodes are then aggregated, forming an optimized and coherent context for the LLM
The Selection Threshold ($S$) governs the scope of data examined, controlling the breadth of information explored. In contrast, the Delta Threshold ($\Delta$) governs the extent of information examined, managing the depth of the search.
Detailed pseudocode for the HIRO Querying Mechanism discussed is available in Algorithm 1 and Algorithm 2.
Experiments
Datasets Used: We measured the performance of HIRO paired with RAPTOR across two question-answering datasets - NarrativeQA and QuALITY.
NarrativeQA is an English-language dataset, that consists of stories, which are books and movie scripts, with human written questions and answers based solely on human-generated abstractive summaries. It tests comprehension with 1567 stories and 46,765 questions (Kočiský et al. 2018), focusing on narratives of 50,000-100,000 words. This dataset tests systems’ ability to prune irrelevant details for key story elements, showcasing querying optimization for narrative understanding and information retrieval. We measure performance on this dataset using the standard ROUGE, BLEU-1, BLEU-4, METEOR.
The QuALITY dataset (Question Answering with Long Input Text, Yes!) is designed for multiple-choice question answering (QA) tasks involving lengthy English texts, ranging from 2000 to 8000 tokens per article. It comprises a total of 6,737 questions, with a challenging subset, QuALITY-HARD, containing 3,360 particularly difficult questions (Pang et al. 2021). This dataset aims to evaluate the ability of QA models to process and extract relevant information from extensive text passages. Performance is measured using the standard F1 score and Accuracy, highlighting the effectiveness of various pruning methods in identifying crucial information within long documents.
Controlled Baseline Comparisons
To establish baseline performance, we utilize RAPTOR with SBERT (Reimers and Gurevych 2019) as the embedding model and GPT 3.5 Turbo (Brown et al. 2020) as the reader/language model. This configuration represents the state-of-the-art approach in retrieval-augmented generation with RAPTOR. We compare the performance of RAPTOR coupled with SBERT and GPT 3.5 Turbo with and without HIRO querying implemented. RAPTOR conventionally selects the top-$k$ nodes after flattening the tree (Collapsed Tree Querying), and we analyze the difference in performance between the conventional approach and HIRO.
Additionally, we provide insights into the token length of the final context returned to the LLM after querying, offering a deeper understanding of the tradeoff between computational resources and performance gains.
Furthermore, the performance of LLMs degrades non-linearly with prompt token length (Levy, Jacoby, and Goldberg 2024), we also present normalized performance scores for conventional querying and HIRO querying, adjusted by the logarithm of context length (in tokens) under efficiency analysis.
Experimental Setup
Experimental Results
Model Performance on Sequence Generation Tasks: Table 1 shows the average context length in tokens returned by different querying mechanisms with the RAPTOR model on the NarrativeQA dataset. The HIRO querying mechanism returned shorter contexts on average compared to tree traversal and collapsed tree querying.
Model Performance on NarrativeQA Dataset: Table 2 presents absolute performance comparison of RAPTOR with collapsed tree querying vs HIRO querying on the NarrativeQA dataset using SBERT and GPT-3.5 Turbo. HIRO querying outperformed conventional querying across all metrics. Table 3 shows the efficiency-adjusted performance scores, calculated by adjusting the absolute scores by the logarithm of context length. HIRO querying demonstrated superior efficiency-adjusted performance over conventional querying on all metrics.
Figure 5 visualizes the absolute impact of HIRO querying, showing the performance gain over collapsed tree querying for ROUGE, BLEU-1, BLEU-4, and METEOR metrics. Figure 6 depicts the efficiency analysis, illustrating that HIRO querying improves both absolute performance as well as efficiency compared to collapsed tree querying.
Model Performance on the QuALITY Dataset: Table 4 compares the absolute accuracy, precision, recall and F1 score of RAPTOR with HIRO querying vs collapsed tree querying on the QuALITY dataset. HIRO querying showed slightly lower accuracy and F1 score, compared to collapsed tree querying.
Conclusion
This research introduces Hierarchical Information Retrieval Optimization (HIRO), a novel technique that enhances Retrieval-Augmented Generation (RAG) systems using hierarchical document structures. By integrating Depth-First Search (DFS)-based recursive similarity thresholding and branch pruning, HIRO effectively manages the context provided to Large Language Models (LLMs), addressing the prevalent challenges of information overload. The approach not only improves the quality and relevance of responses but also demonstrates increased efficiency in data retrieval, reducing computational load and enhancing overall model performance.
Reproducibility Statement
Language Models used for QA and Summarization
OpenAI GPT 3.5 Turbo (Levy, Jacoby, and Goldberg 2024)https://platform.openai.com/docs/models/gpt-3-5-turbo
Datasets Used
NarrativeQA (Kočiský et al. 2018)https://github.com/nyu-mll/quality
QuALITY (Pang et al. 2021)https://github.com/google-deepmind/narrativeqa
Source Code
RAPTOR + HIRO Querying Mechanismhttps://github.com/krishgoel/hiro
Original RAPTOR Source Codehttps://github.com/parthsarthi03/raptor
Appendix A Appendix
Pseudocode for Existing Tree Querying Mechanisms
The following section outlines two distinct conventional querying methods used for querying trees - Tree Traversal Querying in Algorithm 3 and Collapsed Tree Querying in Algorithm 4.
Algorithm 3 Tree Traversal Querying
Input: Tree $tree$, Query $query$, Integer $k$
Output: Top $k$ nodes
[…]
2:  for layer in range($tree$.num_layers) do
3:     $top\_k\leftarrow[]$
4:     for node in $S_{\text{current}}$ do
5:        $score\leftarrow\text{similarity}(query,node)$
6:        $top\_k.\text{append}(node,score)$
7:     end for
8:     $S_{\text{layer}}\leftarrow\text{sorted}(top\_k)[:k].\text{nodes}$
9:     $S_{\text{current}}\leftarrow S_{\text{layer}}$
10:  end for
11:  return $\bigcup_{i=0}^{k}S_{i}$
Algorithm 4 Collapsed Tree Querying
Input: Tree $tree$, Query $query$, Integer $k$, Integer $max\_tokens$
Output: Top nodes within token limit
1:  $tree\leftarrow\text{flatten}(tree)$
2:  $top\_nodes\leftarrow[]$
3:  for node in $tree$ do
4:     $top\_nodes.\text{append}((node,\text{similarity}(query,node)))$
5:  end for
6:  $top\_nodes\leftarrow\text{sorted}(top\_nodes)$
7:  $result\leftarrow[]$
8:  $total\_tokens\leftarrow 0$
9:  for node in $top\_nodes$ do
10:     if $total\_tokens+node\text{.token\_size}<max\_tokens$ then
11:        $result.\text{append}(node)$
12:     end if
[…]
14:  end for
15:  return $result$
Bayesian Optimization for Fine-Tuning of Hyperparameters
@@ Appendix A Appendix
We utilize Bayesian Optimization to fine-tune the hyperparameters Selection Threshold ($S$) and Delta Threshold ($\Delta$) for various datasets. Bayesian Optimization is chosen for its proficiency in optimizing complex, multi-metric objective functions (Snoek, Larochelle, and Adams 2012). Bayesian Optimization utilizes a probabilistic model to estimate the objective function landscape and an acquisition function to efficiently explore this landscape by focusing on areas likely to offer improvements.

### v2

@@ 4.4. Experimental Results
RAPTOR with HIRO Querying 2253.345865
RAPTOR with Collapsed Tree Querying 1864.293445
Table 2. Comparison of Time Complexities of Hierarchical Querying Algorithms
Algorithm Time Complexity
HIRO $O(n)$
Tree Traversal $O(n\log m)$
Collapsed Tree $O(n\log n)$
Table 3. NarrativeQA Dataset: Absolute Performance Metrics and Efficiency-Adjusted Scores. Efficiency scores are adjusted using the logarithm of context length (in tokens) to account for loss in performance with increasing context length.
Model Score Type ROUGE-L F1 BLEU-1 BLEU-4 METEOR
RAPTOR with Collapsed Tree Querying Absolute Performance 0.096542025 0.165350602 0.084131409 0.125686168
RAPTOR with HIRO Querying Absolute Performance 0.120835776 0.173288844 0.089863157 0.089863157
RAPTOR with Collapsed Tree Querying Efficiency Adjusted 0.012819892 0.021957038 0.011171877 0.016689966
RAPTOR with HIRO Querying Efficiency Adjusted 0.015651944 0.022446228 0.011640039 0.017361359
Table 4. Absolute Performance Metrics of Collapsed Tree and HIRO Querying on the QuALITY Dataset.
@@
Model Performance on NarrativeQA Dataset: Table 2 presents absolute performance and efficiency adjusted metrics comparison of RAPTOR with collapsed tree querying vs HIRO querying on the NarrativeQA dataset using SBERT and GPT-3.5 Turbo. HIRO querying outperformed conventional querying across all metrics.
Model Performance on the QuALITY Dataset: Table 3 compares the absolute accuracy, precision, recall and F1 score of RAPTOR with HIRO querying against Collapsed Tree querying on the QuALITY dataset. HIRO querying showed slightly lower accuracy and F1 score, compared to collapsed tree querying.
(a) RAPTOR with HIRO Querying
(b) RAPTOR with Collapsed Tree Querying
@@
Figure 3. Confusion Matrices for RAPTOR with (a) HIRO Querying and (b) Collapsed Tree Querying on the QuALITY dataset
[…]
Confusion Matrices for RAPTOR with (a) HIRO Querying and (b) Collapsed Tree Querying on the QuALITY dataset
Time Complexity Considerations: Table 4 details the time complexities for each hierarchical querying algorithm, highlighting the efficiency of HIRO querying with a linear time complexity $O(n)$, compared to the logarithmic complexities of other methods. This efficiency is crucial for large-scale applications where processing speed is paramount.
For a more detailed analysis of space and time complexity, please refer to Appendix A.
5. Conclusion
This research introduces Hierarchical Information Retrieval Optimization (HIRO), a novel approach that significantly enhances Retrieval-Augmented Generation (RAG) systems. By leveraging hierarchical document structures, HIRO optimizes context management for Large Language Models through recursive similarity thresholding and branch pruning, achieving a more efficient time complexity of O($n$). This method effectively balances context length, avoiding information overload while ensuring sufficient detail. HIRO demonstrates superior performance on the NarrativeQA dataset, surpassing traditional methods in key metrics. Although its performance on the QuALITY dataset shows slightly lower accuracy, it remains competitive. Notably, HIRO’s optimization of context size results in substantial computational efficiency gains, reducing processing costs and paving the way for more sophisticated and resource-efficient AI applications.
6. Reproducibility Statement
Language Model used for QA and Summarization: OpenAI GPT 3.5 Turbo (Brown et al., 2020) (https://platform.openai.com/docs/models/gpt-3-5-turbo)
Datasets Used: NarrativeQA (Kočiský et al., 2018) (https://github.com/google-deepmind/narrativeqa), QuALITY (Pang et al., 2021) (https://github.com/nyu-mll/quality)
RAPTOR with HIRO Querying Mechanism (Source Code): https://github.com/krishgoel/hiro
RAPTOR (Original Source Code): https://github.com/parthsarthi03/raptor
Appendix A Complexity Analysis
In this section, we examine the computational and space complexities of three hierarchical retrieval algorithms—Tree Traversal Querying, Collapsed Tree Querying, and HIRO Hierarchical Querying.
Tree Traversal Querying: This method involves sorting nodes within each layer to select the top $k$ nodes based on similarity. The sorting operation yields a time complexity of $O(m\log m)$, where $m$ is the number of nodes in the largest layer. Consequently, the overall time complexity is $O(n\log m)$, with $n$ representing the total number of nodes. The space complexity is $O(n)$ due to the need to store the hierarchical structure.
Collapsed Tree Querying: This approach flattens the hierarchical structure into a list and sorts the nodes, leading to a time complexity of $O(n\log n)$. The space complexity for storing the hierarchy is also $O(n)$. While straightforward, this method may become less efficient with very large datasets.
HIRO Hierarchical Querying: HIRO employs a dynamic traversal strategy with adaptive thresholds, based on node similarity and parent context, achieving a linear time complexity of $O(n)$ assuming constant times for similarity calculations and threshold adjustments. The space requirement includes $O(n)$ for storing the hierarchy and $O(n\cdot d)$ for node embeddings, where $d$ is the dimensionality of the embeddings. Additional structures also require $O(n)$ space, ensuring overall efficiency.
Figure 4 illustrates the time complexities of these three hierarchical retrieval algorithms.
Figure 4. Comparison of Time Complexities for the Three Hierarchical Retrieval Algorithms.
[…]
Comparison of Time Complexities for the Three Hierarchical Retrieval Algorithms.
Figure 5. Tree Traversal Querying Mechanism. The top-$k$ root nodes are selected based on their similarity to the query. In the next layer, the children of these nodes are considered, and the top-$k$ nodes are again selected based on similarity. This process continues until reaching the leaf nodes.
[…]
The top-$k$ root nodes are selected based on their similarity to the query. In the next layer, the children of these nodes are considered, and the top-$k$ nodes are again selected based on similarity. This process continues until reaching the leaf nodes.
Figure 6. Collapsed Tree Querying Mechanism. The tree is flattened into a single layer, and the top-$k$ nodes are directly selected based on their similarity to the query embedding. The returned context consists of the most relevant nodes from this single-layer selection.
[…]
The tree is flattened into a single layer, and the top-$k$ nodes are directly selected based on their similarity to the query embedding. The returned context consists of the most relevant nodes from this single-layer selection.
Appendix B Production-Ready Pseudocode for HIRO Querying
This revised version of the HIRO Querying pseudocode is designed for optimal integration into production environments. It replaces potentially complex while loops with more manageable for loops, ensuring both clarity and ease of maintenance. Importantly, it integrates two algorithms: HIRO Querying (Algorithm 3) and its auxiliary Evaluate Children (Algorithm 4), which is invoked within the HIRO algorithm. This configuration not only adheres to production standards but also facilitates easier debugging and scalability in operational systems.
While we utilize cosine similarity for node comparisons, other metrics like Euclidean distance, or Manhattan Distance can also be employed to measure vector distances.
@@ Appendix B Production-Ready Pseudocode for HIRO Querying
10:           $local\_context.\text{extend}(\text{evaluateChildren}(q,[node],S,\Delta))$
@@
Appendix C Bayesian Optimization for Fine-Tuning of Hyperparameters
@@
We utilize Bayesian Optimization to fine-tune the hyperparameters Selection Threshold ($S$) and Delta Threshold ($\Delta$) for various datasets. Bayesian Optimization is chosen for its proficiency in optimizing complex, multi-metric objective functions (Snoek et al., 2012). Bayesian Optimization utilizes a probabilistic model to estimate the objective function landscape and an acquisition function to efficiently explore this landscape by focusing on areas likely to offer improvements.